# Distributionally Robust Receive Beamforming

### v2

@@ Abstract
This article investigates signal estimation in wireless transmission (i.e., receive beamforming) from the perspective of statistical machine learning, where the transmit signals may be from an integrated sensing and communication system; that is, 1) signals may be not only discrete constellation points but also arbitrary complex values; 2) signals may be spatially correlated. Particular attention is paid to handling various uncertainties such as the uncertainty of the transmit signal covariance, the uncertainty of the channel matrix, the uncertainty of the channel noise covariance, the existence of channel impulse noises, and the limited sample size of pilots. To proceed, a distributionally robust machine learning framework that is insensitive to the above uncertainties is proposed, which reveals that channel estimation is not a necessary operation. For optimal linear estimation, the proposed framework includes several existing beamformers as special cases such as diagonal loading and eigenvalue thresholding. For optimal nonlinear estimation, estimators are limited in reproducing kernel Hilbert spaces and neural network function spaces, and corresponding uncertainty-aware solutions (e.g., kernelized diagonal loading) are derived. In addition, we prove that the ridge and kernel ridge regression methods in machine learning are distributionally robust against diagonal perturbation in feature covariance.
@@
Wireless Transmission, Smart Antenna, Machine Learning, Robust Estimation, Robust Beamforming, Distributional Uncertainty, Channel Uncertainty, Limited Pilot.
@@ I Introduction
In wireless transmission, detection and estimation of transmitted signals is of high importance, and beamforming at array receivers serves as a key signal-processing technique to suppress interference and environmental noises. The earliest beamforming solutions rely on the use of phase shifters (e.g., phased arrays) to steer and shape wave lobes, while advanced beamforming methods allow the employment of digital signal processing units, which introduce additional structural freedom (e.g., fully digital, hybrid, nonlinear, wideband) in beamformer design and significant performance improvement in signal recovery [1, 2, 3].
@@
In traditional communication systems, transmitted signals are discrete points from constellations. Therefore, signal recovery, commonly referred to as signal detection, can be cast into a classification problem from the perspective of statistical machine learning, and the number of candidate classes is determined by the number of points in the employed constellation. Research in this stream includes, e.g., [4, 5, 6, 7, 8, 9] as well as references therein, and the performance measure for signal detection is usually the misclassification rate (i.e., symbol error rate); representative algorithms encompass the maximum likelihood detector, the sphere decoding, etc. In another research stream, the signal recovery performance is evaluated using mean-squared errors (cf., signal-to-interference-plus-noise ratio), and the resultant signal recovery problem is commonly known as signal estimation, which can be considered as a regression problem from the perspective of statistical machine learning. By comparing the estimated symbols with the constellation points afterward, the detection of discrete symbols can be realized. For this case, till now, typical beamforming solutions include zero-forcing receivers, Wiener receivers (i.e., linear minimum mean-squared error receivers), Capon receivers (i.e., minimum variance distortionless response receivers), and nonlinear receivers such as neural-network receivers [10, 11, 12]. On the basis of these canonical approaches, variants such as robust beamformers working against the limited size of pilot samples and the uncertainty in steering vectors [13, 14, 15, 16, 17, 18] have also been intensively reported; among these robust solutions, the diagonal loading method [19], [14, Eq. (11)] and the eigenvalue thresholding method [20], [14, Eq. (12)], etc., are popular due to their excellent balance in practical performance and technical simplicity.
@@
Different from traditional paradigms, in emerging communication systems, e.g., integrated sensing and communication (ISAC) systems, transmitted signals may be arbitrary complex-valued symbols and spatially correlated [21, 22, 23]. As a result, mean-squared error is a preferred performance measure to investigate the receive beamforming and estimation problem of wireless signals, which is, therefore, the focus of this article.
@@
Although a large body of problems have been attacked in the area, the following signal-processing problems of beamforming and estimation in wireless transmission remain unsolved.
@@
    What is the relation between the signal-model-based approaches (e.g., Wiener and Capon receivers) and the data-driven approaches (e.g., deep-learning receivers)? In other words, how can we build a mathematically unified modeling framework to interpret all the existing digital receive beamformers?
@@
    In addition to the limited pilot size and the uncertainty in steering vectors, there exist other uncertainties in the signal model: the uncertainty in the transmit signal covariance, the uncertainty of the communication channel matrix, the uncertainty of the channel noise covariance, and the presence of channel impulse noises (i.e., outliers). Therefore, how can we handle all these types of uncertainties in a unified solution framework?
@@
    Existing literature mainly studied the robustness theory of linear beamformers against limited pilot size and the uncertainty in steering vectors [13, 14, 15, 16, 17, 18]. However, how can we develop the theory of robust nonlinear beamformers against all the aforementioned uncertainties?
To this end, this article designs a unified modeling and solution framework for receive beamforming of wireless signals, in consideration of the scarcity of the pilot data and the different uncertainties in the signal model.
@@ I-A Contributions
    We build a fundamentally theoretical framework for receive beamforming from the perspective of statistical machine learning. In addition to the linear estimation methods, nonlinear approaches (i.e., nonlinear beamforming) are also discussed in reproducing kernel Hilbert spaces and neural network function spaces. In particular, we reveal that channel estimation is not a necessary operation in receive beamforming. For details, see Subsection III-A.
@@
    The presented framework is particularly developed from the perspective of distributional robustness which can therefore combat the limited size of pilot data and several types of uncertainties in the wireless signal model such as the uncertainty in the transmit power matrix, the uncertainty in the communication channel matrix, the existence of channel impulse noises (i.e., outliers), the uncertainty in the covariance matrix of channel noises, etc. For details, see Subsection III-B, and the technical developments in Sections IV and V.
@@
    The distributionally robust receive beamforming and signal estimation problems across multiple frames, where channel conditions may change, are also investigated. For details, see Subsections IV-C and V-A2.
@@ II Preliminaries
We review two popular structured representation methods of nonlinear functions $\bm{\phi}:\mathbb{R}^{N}\to\mathbb{R}^{M}$. More details can be seen in Appendix A of the online supplementary materials.
@@ II-B Neural Networks
Neural networks (NN) are another powerful tool to represent (i.e., approximate) nonlinear functions. A neural network function space (NNFS) $\mathcal{K}$ characterizes (or parameterizes) a set of multi-input multi-output functions. Typical choices are multi-layer feed-forward neural networks, recurrent neural networks, etc. For beamforming and estimation of wireless signals, the multi-layer feed-forward neural networks are standard [10, 11, 12]. Suppose that we have $R-1$ hidden layers (so in total $R+1$ layers including one input layer and one output layer) and each layer $r=0,1,\ldots,R$ contains $T_{r}$ neurons. To represent a function $\bm{\phi}:\mathbb{R}^{N}\to\mathbb{R}^{M}$, for the input layer $r=0$ and output layer $r=R$, we have $T_{0}=N$ and $T_{R}=M$, respectively. Let the output of the $r^{\text{th}}$ layer be $\bm{y}_{r}\in\mathbb{R}^{T_{r}}$. For every layer $r$, we have $\bm{y}_{r}=\bm{\sigma}_{r}(\bm{W}^{\circ}_{r}\cdot\bm{y}_{r-1}+\bm{b}_{r})$ where $\bm{W}^{\circ}_{r}\in\mathbb{R}^{T_{r}\times T_{r-1}}$ is the weight matrix, $\bm{b}_{r}\in\mathbb{R}^{T_{r}}$ is the bias vector, and the multi-output function $\bm{\sigma}_{r}$ is the activation function which is entry-wise identical. Hence, every function $\bm{\phi}:\mathbb{R}^{N}\to\mathbb{R}^{M}$ in a NNFS can be recursively expressed as [25, Chap. 5], [26]
@@ III Problem Formulation
Consider a narrow-band wireless signal model
@@
where $\mathbf{x}\in\mathbb{C}^{N}$ is the received signal, $\mathbf{s}\in\mathbb{C}^{M}$ is the transmitted signal, $\bm{H}\in\mathbb{C}^{N\times M}$ is the channel matrix, and $\mathbf{v}\in\mathbb{C}^{N}$ is the zero-mean channel noise. The precoding operation (if exists) is integrated in $\bm{H}$. The transmitted symbols $\mathbf{s}$ have zero means, which may be not only discrete values from constellations such as quadrature amplitude modulation but also arbitrary symbols such as integrated sensing and communication (ISAC) signals. We consider $L$ pilots $\bm{S}\coloneqq(\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L})$ in each frame, and the corresponding received symbols are $\bm{X}\coloneqq(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{L})$ under the noise $(\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{L})$. We suppose that $\bm{R}_{s}\coloneqq\mathbb{E}\mathbf{s}\mathbf{s}^{\mathsf{H}}$ and $\bm{R}_{v}\coloneqq\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{H}}$ may not be identity or diagonal matrices: i.e., the components of $\mathbf{s}$ can be correlated (e.g., in ISAC), so can be these of $\mathbf{v}$. Consider the real-space representation of the signal model (5) by stacking the real and imaginary components:
@@
where $\underline{\mathbf{x}}\in\mathbb{R}^{2N}$, $\underline{\underline{\bm{H}}}\in\mathbb{R}^{2N\times 2M}$, $\underline{\mathbf{s}}\in\mathbb{R}^{2M}$, and $\underline{\mathbf{v}}\in\mathbb{R}^{2N}$. The expressions of $\bm{R}_{\underline{x}}\coloneqq\mathbb{E}{\underline{\mathbf{x}}\underline{\mathbf{x}}^{\mathsf{T}}}$, $\bm{R}_{\underline{s}}\coloneqq\mathbb{E}{\underline{\mathbf{s}}\underline{\mathbf{s}}^{\mathsf{T}}}$, $\bm{R}_{\underline{x}\underline{s}}\coloneqq\mathbb{E}{\underline{\mathbf{x}}\underline{\mathbf{s}}^{\mathsf{T}}}$, and $\bm{R}_{\underline{v}}\coloneqq\mathbb{E}{\underline{\mathbf{v}}\underline{\mathbf{v}}^{\mathsf{T}}}$ can be readily obtained; see Appendix B of the online supplementary materials. Signal estimation in real spaces can be technically simpler than that in complex spaces.
@@ III-A Optimal Estimation
III-A1 Optimal Nonlinear Estimation (Nonlinear Beamforming)
To recover $\mathbf{s}$ using $\mathbf{x}$, we consider an estimator $\hat{\mathbf{s}}\coloneqq\bm{\phi}(\mathbf{x})$ at the receiver where $\bm{\phi}:\mathbb{C}^{N}\to\mathbb{C}^{M}$ is a Borel-measurable function; note that $\bm{\phi}(\mathbf{x})$ may be nonlinear in general. The signal estimation problem at the receiver can be written as a statistical machine-learning problem under the joint data distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ of $(\mathbf{x},\mathbf{s})$, that is,
@@ III-A1 Optimal Nonlinear Estimation (Nonlinear Beamforming)
Therefore, we have $\mathbf{s}=\hat{\mathbf{s}}+\mathbf{e}$, that is,
$\mathbf{s}=\bm{\phi}(\mathbf{x})+\mathbf{e}$ (9)
where $\mathbf{e}$ denotes the signal’s estimation error; (9) implies that, in practice where $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is unknown, the data-driven estimation problem of $\mathbf{s}$ given $\mathbf{x}$ can be seen as a nonlinear regression problem of the collected data pairs $\{(\bm{x}_{1},\bm{s}_{1}),(\bm{x}_{2},\bm{s}_{2}),\ldots,(\bm{x}_{L},\bm{s}_{L})\}$. Usually, it is technically complicated to find the optimal $\bm{\phi}(\cdot)$ from the whole space $\mathcal{B}$ of Borel-measurable functions. Therefore, in practice, we may find the optimal approximation of $\bm{\phi}(\cdot)$ in an RKHS $\mathcal{H}$ or a NNFS $\mathcal{K}$; note that $\mathcal{H}$ and $\mathcal{K}$ are two subspaces of $\mathcal{B}$. However, both $\mathcal{H}$ and $\mathcal{K}$ are sufficiently rich because they can be dense in the space of all continuous bounded functions.
III-A2 Optimal Linear Estimation (Beamforming)
When $\mathbf{x}$ and $\mathbf{s}$ are jointly Gaussian (e.g., when $\mathbf{s}$ and $\mathbf{v}$ are jointly Gaussian), the optimal estimator $\bm{\phi}$ is linear in $\mathbf{x}$:
@@ III-A2 Optimal Linear Estimation (Beamforming)
where $\bm{W}\in\mathbb{C}^{M\times N}$ is called a smart antenna or a receive beamformer. In this linear case, (7) reduces to the usual Wiener–Hopf beamforming problem
@@
which is known as the Wiener beamformer. With an additional constraint $\bm{W}\bm{H}=\bm{I}_{M}$ (i.e., distortionless response), (12) gives the Capon beamformer. Both the Wiener beamformer and the Capon beamformer maximize the output signal–to–interference-plus-noise ratio (SINR); hence, both are optimal in the sense of maximum output SINR. No matter whether $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is Gaussian or not, (11) or (12) gives the optimal linear beamformer in the sense of maximizing the output SINR among all linear beamformers.

### v3

# Distributionally Robust Receive Combining

@@ Abstract
This article investigates signal estimation in wireless transmission (i.e., receive combining) from the perspective of statistical machine learning, where the transmit signals may be from an integrated sensing and communication system; that is, 1) signals may be not only discrete constellation points but also arbitrary complex values; 2) signals may be spatially correlated. Particular attention is paid to handling various uncertainties such as the uncertainty of the transmit signal covariance, the uncertainty of the channel matrix, the uncertainty of the channel noise covariance, the existence of channel impulse noises, the non-ideality of the power amplifiers, and the limited sample size of pilots. To proceed, a distributionally robust receive combining framework that is insensitive to the above uncertainties is proposed, which reveals that channel estimation is not a necessary operation. For optimal linear estimation, the proposed framework includes several existing combiners as special cases such as diagonal loading and eigenvalue thresholding. For optimal nonlinear estimation, estimators are limited in reproducing kernel Hilbert spaces and neural network function spaces, and corresponding uncertainty-aware solutions (e.g., kernelized diagonal loading) are derived. In addition, we prove that the ridge and kernel ridge regression methods in machine learning are distributionally robust against diagonal perturbation in feature covariance.
@@
Wireless Transmission, Smart Antenna, Machine Learning, Robust Estimation, Robust Combining, Distributional Uncertainty, Channel Uncertainty, Limited Pilot.
@@ I Introduction
In wireless transmission, detection and estimation of transmitted signals is of high importance, and combining at array receivers serves as a key signal-processing technique to suppress interference and environmental noises. The earliest beamforming solutions rely on the use of phase shifters (e.g., phased arrays) to steer and shape wave lobes, while advanced combining methods allow the employment of digital signal processing units, which introduce additional structural freedom (e.g., fully digital, hybrid, nonlinear, wideband) in combiner design and significant performance improvement in signal recovery [1, 2, 3].
@@
In traditional communication systems, transmitted signals are discrete points from constellations. Therefore, signal recovery, commonly referred to as signal detection, can be cast into a classification problem from the perspective of statistical machine learning, and the number of candidate classes is determined by the number of points in the employed constellation. Research in this stream includes, e.g., [4, 5, 6, 7, 8, 9] as well as references therein, and the performance measure for signal detection is usually the misclassification rate (i.e., symbol error rate); representative algorithms encompass the maximum likelihood detector, the sphere decoding, etc. In another research stream, the signal recovery performance is evaluated using mean-squared errors (cf., signal-to-interference-plus-noise ratio), and the resultant signal recovery problem is commonly known as signal estimation, which can be considered as a regression problem from the perspective of statistical machine learning. By comparing the estimated symbols with the constellation points afterward, the detection of discrete symbols can be realized. For this case, till now, typical combining solutions include zero-forcing receivers, Wiener receivers (i.e., linear minimum mean-squared error receivers), Capon receivers (i.e., minimum variance distortionless response receivers), and nonlinear receivers such as neural-network receivers [10, 11, 12]. On the basis of these canonical approaches, variants such as robust beamformers working against the limited size of pilot samples and the uncertainty in steering vectors [13, 14, 15, 16, 17, 18] have also been intensively reported; among these robust solutions, the diagonal loading method [19], [14, Eq. (11)] and the eigenvalue thresholding method [20], [14, Eq. (12)] are popular due to their excellent balance between practical performance and technical simplicity.
@@
Different from traditional paradigms, in emerging communication systems, e.g., integrated sensing and communication (ISAC) systems, transmitted signals may be arbitrary complex values and spatially correlated [21, 22, 23]. As a result, mean-squared error is a preferred performance measure to investigate the receive combining and estimation problem of wireless signals, which is, therefore, the focus of this article.
@@
Although a large body of problems have been attacked in the area, the following signal-processing problems of combining and estimation in wireless transmission remain unsolved.
@@
    What is the relation between the signal-model-based approaches (e.g., Wiener and Capon receivers) and the data-driven approaches (e.g., deep-learning receivers)? In other words, how can we build a mathematically unified modeling framework to interpret all the existing digital receive combiners?
@@
    In addition to the limited pilot size and the uncertainty in steering vectors, there exist other uncertainties in the signal model: the uncertainty of the transmit signal covariance, the uncertainty of the communication channel matrix, the uncertainty of the channel noise covariance, the presence of channel impulse noises (i.e., outliers), and the non-ideality of the power amplifiers. Therefore, how can we handle all these types of uncertainties in a unified solution framework?
@@
    Existing literature mainly studied the robustness theory of linear beamformers against limited pilot size and the uncertainty in steering vectors [13, 14, 15, 16, 17, 18]. However, how can we develop the theory of robust nonlinear combiners against all the aforementioned uncertainties?
To this end, this article designs a unified modeling and solution framework for receive combining of wireless signals, in consideration of the scarcity of the pilot data and the different uncertainties in the signal model.
@@ I-A Contributions
    We build a fundamentally theoretical framework for receive combining from the perspective of statistical machine learning. In addition to the linear estimation methods, nonlinear approaches (i.e., nonlinear combining) are also discussed in reproducing kernel Hilbert spaces and neural network function spaces. In particular, we reveal that channel estimation is not a necessary operation in receive combining. For details, see Subsection III-A.
@@
    The presented framework is particularly developed from the perspective of distributional robustness which can therefore combat the limited size of pilot data and several types of uncertainties in the wireless signal model such as the uncertainty in the transmit power matrix, the uncertainty in the communication channel matrix, the existence of channel impulse noises (i.e., outliers), the uncertainty in the covariance matrix of channel noises, the non-ideality of the power amplifiers, etc. For details, see Subsection III-B, and the technical developments in Sections IV and V.
@@
    The distributionally robust receive combining and signal estimation problems across multiple frames, where channel conditions may change, are also investigated. For details, see Subsections IV-C and V-A2.
@@ II Preliminaries
We review two popular structured representation methods of nonlinear functions $\bm{\phi}:\mathbb{R}^{N}\to\mathbb{R}^{M}$. More details can be seen in Appendix A.
@@ II-B Neural Networks
Neural networks (NN) are another powerful tool to represent (i.e., approximate) nonlinear functions. A neural network function space (NNFS) $\mathcal{K}$ characterizes (or parameterizes) a set of multi-input multi-output functions. Typical choices are multi-layer feed-forward neural networks, recurrent neural networks, etc. For combining and estimation of wireless signals, the multi-layer feed-forward neural networks are standard [10, 11, 12]. Suppose that we have $R-1$ hidden layers (so in total $R+1$ layers including one input layer and one output layer) and each layer $r=0,1,\ldots,R$ contains $T_{r}$ neurons. To represent a function $\bm{\phi}:\mathbb{R}^{N}\to\mathbb{R}^{M}$, for the input layer $r=0$ and output layer $r=R$, we have $T_{0}=N$ and $T_{R}=M$, respectively. Let the output of the $r^{\text{th}}$ layer be $\bm{y}_{r}\in\mathbb{R}^{T_{r}}$. For every layer $r$, we have $\bm{y}_{r}=\bm{\sigma}_{r}(\bm{W}^{\circ}_{r}\cdot\bm{y}_{r-1}+\bm{b}_{r})$ where $\bm{W}^{\circ}_{r}\in\mathbb{R}^{T_{r}\times T_{r-1}}$ is the weight matrix, $\bm{b}_{r}\in\mathbb{R}^{T_{r}}$ is the bias vector, and the multi-output function $\bm{\sigma}_{r}$ is the activation function which is entry-wise identical. Hence, every function $\bm{\phi}:\mathbb{R}^{N}\to\mathbb{R}^{M}$ in a NNFS can be recursively expressed as [25, Chap. 5], [26]
@@ III Problem Formulation
Consider a narrow-band wireless signal transmission model
@@
where $\mathbf{x}\in\mathbb{C}^{N}$ is the received signal, $\mathbf{s}\in\mathbb{C}^{M}$ is the transmitted signal, $\bm{H}\in\mathbb{C}^{N\times M}$ is the channel matrix, and $\mathbf{v}\in\mathbb{C}^{N}$ is the zero-mean channel noise. The precoding operation (if exists) is integrated in $\bm{H}$. The transmitted symbols $\mathbf{s}$ have zero means, which may be not only discrete symbols from constellations such as quadrature amplitude modulation but also arbitrary values such as integrated sensing and communication signals. We consider $L$ pilots $\mathbf{S}\coloneqq(\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{L})$ in each frame, and the corresponding received symbols are $\mathbf{X}\coloneqq(\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{L})$ under the noise $(\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{L})$. We suppose that $\bm{R}_{s}\coloneqq\mathbb{E}\mathbf{s}\mathbf{s}^{\mathsf{H}}$ and $\bm{R}_{v}\coloneqq\mathbb{E}\mathbf{v}\mathbf{v}^{\mathsf{H}}$ may not be identity or diagonal matrices: i.e., the components of $\mathbf{s}$ can be correlated (e.g., in ISAC), so can be these of $\mathbf{v}$. Consider the real-space representation of the signal model (5) by stacking the real and imaginary components:
@@
where $\underline{\mathbf{x}}\in\mathbb{R}^{2N}$, $\underline{\underline{\bm{H}}}\in\mathbb{R}^{2N\times 2M}$, $\underline{\mathbf{s}}\in\mathbb{R}^{2M}$, and $\underline{\mathbf{v}}\in\mathbb{R}^{2N}$. The expressions of $\bm{R}_{\underline{x}}\coloneqq\mathbb{E}{\underline{\mathbf{x}}\underline{\mathbf{x}}^{\mathsf{T}}}$, $\bm{R}_{\underline{s}}\coloneqq\mathbb{E}{\underline{\mathbf{s}}\underline{\mathbf{s}}^{\mathsf{T}}}$, $\bm{R}_{\underline{x}\underline{s}}\coloneqq\mathbb{E}{\underline{\mathbf{x}}\underline{\mathbf{s}}^{\mathsf{T}}}$, and $\bm{R}_{\underline{v}}\coloneqq\mathbb{E}{\underline{\mathbf{v}}\underline{\mathbf{v}}^{\mathsf{T}}}$ can be readily obtained; see Appendix B. In some cases, signal estimation in real spaces can be technically simpler than that in complex spaces.
@@ III-A Optimal Estimation
III-A1 Optimal Nonlinear Estimation (Receive Combining)
To recover $\mathbf{s}$ using $\mathbf{x}$, we consider an estimator $\hat{\mathbf{s}}\coloneqq\bm{\phi}(\mathbf{x})$, called a receive combiner, at the receiver where $\bm{\phi}:\mathbb{C}^{N}\to\mathbb{C}^{M}$ is a Borel-measurable function. Note that $\bm{\phi}(\mathbf{x})$ may be nonlinear in general because the joint distribution of $(\mathbf{x},\mathbf{s})$ is not necessarily Gaussian, for example, when the channel noise $\mathbf{v}$ is non-Gaussian or when the power amplifiers work in non-linear regions. The signal estimation problem at the receiver can be written as a statistical machine-learning problem under the joint data distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ of $(\mathbf{x},\mathbf{s})$, that is,
@@ III-A1 Optimal Nonlinear Estimation (Receive Combining)
Usually, it is computationally complicated to find the optimal $\bm{\phi}(\cdot)$ from the whole space $\mathcal{B}$ of Borel-measurable functions, that is, to compute the conditional mean. Therefore, in practice, we may find the optimal approximation of $\bm{\phi}(\cdot)$ in an RKHS $\mathcal{H}$ or a NNFS $\mathcal{K}$; note that $\mathcal{H}$ and $\mathcal{K}$ are two subspaces of $\mathcal{B}$. However, both $\mathcal{H}$ and $\mathcal{K}$ are sufficiently rich because they can be dense in the space of all continuous bounded functions.
III-A2 Optimal Linear Estimation (Receive Beamforming)
If $\mathbf{x}$ and $\mathbf{s}$ are jointly Gaussian (e.g., when $\mathbf{s}$ and $\mathbf{v}$ are jointly Gaussian), the optimal estimator $\bm{\phi}$ is linear in $\mathbf{x}$:
@@ III-A2 Optimal Linear Estimation (Receive Beamforming)
where $\bm{W}\in\mathbb{C}^{M\times N}$ is called a receive beamformer or a linear receive combiner. In this linear case, (7) reduces to the usual Wiener–Hopf beamforming problem
@@
which is known as the Wiener beamformer. With an additional constraint $\bm{W}\bm{H}=\bm{I}_{M}$ (i.e., distortionless response), (11) gives the Capon beamformer. Both the Wiener beamformer and the Capon beamformer maximize the output signal–to–interference-plus-noise ratio (SINR); hence, both are optimal in the sense of maximum output SINR.
No matter whether $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is Gaussian or not, (10) or (11) identifies the optimal linear estimator in the sense of minimum mean-squared error among all linear estimators.